This problem is actually for ALG 1 but I can't remember how to complete it? Please help?!
PROBLEM:
The second side of a triangular deck is 5 feet longer than the shortest side and a third side that is 5 feet shorter than twice the length of the shortest side. if the perimeter of the deck is 68 feet, what are the lengths of the three sides?
ANSWER:
The shortest side of the deck is...feet?
The second side of the deck is...feet?
The third side of the deck is… feet?
(Simplify all answers)
So that's the ALG 1 word problem. Please help? I'm stumped on how to solve this math problem.

Answer by lwsshak3(11628)

The second side of a triangular deck is 5 feet longer than the shortest side and a third side that is 5 feet shorter than twice the length of the shortest side. if the perimeter of the deck is 68 feet, what are the lengths of the three sides?
***
let x=length of shortest side
x+5=length of 2nd side (5 ft longer than shortest side)
2x-5=length of 3rd side (twice the length of shortest side-5
x+x+5+2x-5=68
4x=68
x=17
x+5=22
2x-5=29
length of shortest side=17
length of 2nd side=22
length of 3rd side=29